OpenStudy (anonymous):

The length of a rectangle exceeds its width by 4 inches. Find the dimensions of the rectangle if its area is 96 square in.

OpenStudy (sidsiddhartha):

what is the area of a rectangle do u know?

OpenStudy (anonymous):

the formula is length x width

OpenStudy (ranga):

Let W be the width. Length = W + 4 Area = L*W = (W+4)*W = 96 Solve for W and then L.